OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 15x - 5y = 30. intercept = (-2, 0); y-intercept = (0, 6) intercept = (0, -6); y-intercept = (2, 0) x-intercept = (2, 0); y-intercept = (0, -6) intercept = (0, 6); y-intercept = (-2, 0)

OpenStudy (anonymous):

x-intercept is when the graph hits the x-axis therefore y=0 set y=0 in the equation and solve for x \[15x-5(0)=30\]\[x=2\]y-interept is when x=0 repeat the process

OpenStudy (johnweldon1993):

What you have is standard form...put it into slope intercept form 15x - 5y = 30. -5y = 30 - 15x y = -6 + 3x y = 3x - 6 now plug 0 in for y...and then x to find both intersects